Archie Co. purchased a framing machine for $53,000 on January 1, 2021. The machine is expected to have a four-year life, with a residual value of $6,000 at the end of four years.

Using the sum-of-the-years'-digits method, depreciation for 2021 and book value at December 31, 2021, would be: (Do not round intermediate calculations.)

Multiple Choice

  • $21,200 and $31,800 respectively.

  • $18,800 and $34,200 respectively.

  • $18,800 and $28,200 respectively.

  • $21,200 and $25,800 respectively.
Answer #1

Sum = 4*(4+1)/2 = 10

Depreciation for year 1 = (53,000-6000)*4/10

= 18,800

Book value = 53,000 - 18,800 = 34,200

Option B is the answer